q #9
FIND THE y-INTERCEPT OF THE LINE 5x-4y=8

Answer by rfer(16322)

y int is whe x=0
5x-4y=8
5(0)-4y=8
-4y=8
divide by -4
-4y/-4=8/-4
y=-2
(0,-2)